OpenStudy (anonymous):

What is square root of 128 describe

OpenStudy (jdoe0001):

 \(\bf 128 = 2 \times 2 \times 2 \times 2 \times 2 \times 2 \times 2 \implies (2^2)^3 \implies (2^3)^2\\ \sqrt{128} \implies \sqrt{ (2^3)^2}\)